OpenStudy (anonymous):

Exponential and logarithmic models help!! The sales S (in thousands of units of a new CD burner after it has been on the market t years are modeled by S(t)=100(1-e^kt). Fifteen thousand units of the new product were sold in the first year. a) Complete the model by solving for k. b) Sketch the graph of the model. c) Use the model to estimate the number of units sold after 5 years.

OpenStudy (amistre64):

s = 100(1 - e^kt) s = 100 - 100e^kt 100-s = 100e^kt (100-s)/100 = e^kt ln((100-s)/100) = ln(e^kt) ln(100-s) - ln(100) = kt t ln(100-s) - t ln(100) = k ln(100-s)^t - ln(100)^t = k Thats the best I can get you for. and I aint none to sure bout it either :)

OpenStudy (anonymous):

Thank you! I got the correct value of k actually I just found out. Do you know if I would just put the whole equation into my graphing calculator then?

OpenStudy (amistre64):

I dont know, couldnt hurt to try. Ive never used one before :)

OpenStudy (anonymous):

Oh alright thank you for your help! :)